|Operation Summary for March 7, 2012 - March 9, 2012|
Wednesday March 7
During the day shift, at 10:17 AM, a RF person reported finding a severe LCW leak on the cavity tuner for MIRF8; they estimated eight hours to complete this repair. Much of the accelerator complex was accessed for maintenance work. Linac personnel established beam to the Linac dump at 3:43 PM.
During the evening shift, a MiniBooNE person reported at 4:09 PM that their air conditioning installation at MI-12 had been completed. Operations reported at 4:43 PM that all of Switchyard (SY), Meson, the Neutrino Area, NuMI, MiniBooNE, and the Transfer Hall power supplies had been turned on. The power supplies were on. A Booster person established beam to the Booster dump at 4:48 PM. Operators began turning on the Main Injector (MI) at 5:45 PM. Operators had the Meson MTest enclosures searched and secured by 5:46 PM. Operations established beam to NuMI and MiniBooNE at 6:19 PM. Operations established beam to the Meson FTBF experiment T-1008 at 6:31 PM. External beamline personnel began beam commissioning for the Neutrino Area SeaQuest beam at 6:41 PM. Operators completed a Booster Injection tune at 8:12 PM. Operations and External beamline personnel reported at 11:57 PM that earlier they had steered beam through SY to the NM3 target.
Thursday March 8
The midnight shift began with Operations sending beam to the NuMI, MiniBooNE, SeaQuest, and T-1008 experiments.
During the midnight shift, Operations disabled beam to SeaQuest around 12:30 AM. Operators loaded protons into the Recycler at 1:15 AM. Operations disabled beam to T-1008 at 2:03 AM. At 2:43 AM, BRF2 tripped off for the first of three times in about 90 minutes; it reset. Operators completed a Booster Injection tune at 3:30 AM. An External beamline person resumed SeaQuest beam commissioning at 7 AM. At 7:35 AM, Neutron Therapy Facility (NTF) personnel prepared to treat one patent later in the morning.
During the day shift, Operations established beam to T-1008 at 10:14 AM. BRF2 tripped off at 10:20 AM; it reset. BRF14, 15, and 16 all tripped off at 10:54 AM; they all reset. At 11:09 AM, BRF14 tripped off for the first of three times in 43 minutes; it reset. A Booster person began a study at 2:01 PM. At 2:41 PM, BRF14 tripped off for the first of three times in eight minutes; it reset. Operations turned off the Linac beam at 3:02 PM to allow Linac technicians and Water Group personnel to replace a valve on the 400 MeV area water skid; they completed the job and Linac resumed sending beam at 3:39 PM.
During the evening shift, Operations disabled beam to SeaQuest and prepared NM3 and NM4 for access at 5:11 PM. Operations disabled T-1008 beam around 10 PM. Operators completed a Booster injection tune at 11:06 PM.
Friday March 9
The midnight shift began with Operations sending beam to NuMI and MiniBooNE.
During the midnight shift, at 5:06 AM, LRF5 tripped off and wouldn’t reset; a Linac person investigated and then called in people to help with repairs. Operations loaded an energy conservation timeline into the Time Line Generator at 5:36 AM. Operations received a Booster high water sump alarm at 7:02 AM. Operators prepared the Neutrino Area enclosure N01 for access at 7:11 AM. Operators prepared Booster for an access at 7:18 AM; FESS personnel found and temporarily repaired a stuck float. At 8 AM, NTF prepared to treat one patient later in the day.
Operations Coordinator (OpsCo) Plans for Today and the Weekend
The OpsCo agrees with the plans by Booster and Muon personnel to conduct studies and for Operations and external beamline personnel to continue with SeaQuest commissioning.
Comments and Suggestions
|last modified 03/09/2012 email Fermilab|